# Mod $\ell$ non-vanishing of self-dual Hecke $L$-values over CM fields and applications

### v1

@@ Abstract.
Let $\lambda$ be a self-dual Hecke character over a CM field $K$. Let $\mathfrak{p}$} be a prime of the maximal totally real subfield $F$ of $K$ and $\Gamma_{{\mathfrak{p}}}$ the Galois group of the maximal anticyclotomic ${\mathbb{Z}}_{p}^{\deg{\mathfrak{p}}}$-extension of $K$ unramified outside $\mathfrak{p}$}. We prove that
@@
Our approach and results complement the prior work initiated by Hida’s ideas on the arithmetic of Hilbert modular Eisenstein series, studied via mod $\ell$ analogue of the André–Oort conjecture. The previous results established the non-vanishing only for a Zariski dense subset of characters $\nu$. Our approach is based on the arithmetic of a CM modular form on a Shimura set, studied via arithmetic of the CM field and Ratner’s ergodicity of unipotent flows.
@@ 1. Introduction
A basic example arises from CM motives or Hecke characters over a CM field. The aim of this paper is to establish mod $\ell$ non-vanishing of central Hecke $L$-values in a $p$-adic self-dual family over a CM field for primes $\ell\neq p$ (see Theorems 1.1 and 1.4). It has applications to CM Iwasawa theory: a rigidity for Katz $p$-adic $L$-function (see Theorem 1.9) and completion of Hsieh’s proof of Eisenstein congruence divisibility towards the CM main conjecture [28] (see Theorem 1.6).
@@ 1.1. Context
Hida’s mod $\ell$ non-vanishing of Hecke $L$-values over a general CM field has found applications to the arithmetic of the CM field (cf. [22, 41]). It is an ingredient in Hsieh’s 2014 proof of Eisenstein congruence divisibility [28] towards the CM Iwasawa main conjecture over a CM field. However, in 2018 a gap was found in Hida’s strategy, specifically in his proof of Zariski density of CM points on a self-product of $\mathcal{X}$}. In a recent paper [25] Hida recovered [^2] For the case $F={\mathbb{Q}}$, see also [6] and [42]. several of the results of [21, 23, 26]. However, if $\deg{\mathfrak{p}}=1$, then the non-vanishing is only proved for infinitely many characters $\nu$ and not all except finitely many $\nu$ as originally claimed. Consequently, Hsieh’s proof [28] of $\ell$-indivisibility of an ancillary Eisenstein series on the unitary group $U(2,1)$, and hence that of the main result of [28] towards the CM Iwasawa main conjecture are incomplete.
In this paper we establish a uniform non-vanishing of self-dual Hecke $L$-values over a general CM field $K$ by a new method. For a self-dual Hecke character $\lambda$ over $K$, we prove that
@@
for all except finitely many finite order characters $\nu$ of $\Gamma_{\mathfrak{p}}$ such that $\varepsilon(\lambda\nu)=+1$ (see Theorem 1.1). We emphasise that this non-vanishing holds independent of the degree of $\mathfrak{p}$}, being the first result of its kind if $\deg{\mathfrak{p}}>1$. We also prove an analogous mod $\ell$ non-vanishing for primes $\ell\neq p$ ordinary with respect to the CM quadratic extension $K/F$ (see Theorem 1.4). The latter leads to completion of Hsieh’s proof of Eisenstein congruence divisibility towards the CM main conjecture (see  Theorem 1.6). The finer non-vanishing also implies a rigidity for zero divisor of anticyclotomic Katz $\mathfrak{p}$}-adic $L$-function and points to new phenomena (see Theorem 1.9).
Our approach is rooted in the arithmetic of a CM modular form on Shimura set arising from a definite quaternion algebra over $F$ (cf. [6]). It relies on the arithmetic of the CM field, the Waldspurger formula and Ratner’s ergodicity of unipotent flows  [46]. To link arithmetic of the Shimura set with that of the CM field, a key is an $\ell$-integral comparison of quaternionic and CM periods (see Theorem 1.12), which maybe of independent interest.
@@ 1.2.1. Set-up
Let $\lambda$ be a self-dual333conjugate symplectic self-dual in the sense of [17] Hecke over $K$, i.e.
@@
The infinity type of $\lambda$ is of the form $\Sigma+\kappa(1-c)$ for $\Sigma$ a CM type of $K$ and $\kappa=\sum_{\sigma\in\Sigma}\kappa_{\sigma}\cdot\sigma\in{\mathbb{Z}}_{\geq 0}[\Sigma]$, i.e.
@@
$\Xi_{\lambda,{\mathfrak{p}}}^{+}=\{\nu\in\Xi_{\lambda}|\varepsilon(\lambda\nu)=+1\},$
@@
The literature on related problems [21, 23, 24, 25, 12, 13] suggests that such a non-vanishing holds for a Zariski dense [^4] Let $\{\gamma_{1},\cdots,\gamma_{\deg{\mathfrak{p}}}\}$ be a set of topological generators of $\Gamma_{\mathfrak{p}}$. Regard the set of continuous characters ${\mathrm{Hom}}(\Gamma_{\mathfrak{p}},\mu_{p^{\infty}})$ as a subset of ${\mathbb{G}}_{m}^{\deg{\mathfrak{p}}}(\overline{{\mathbb{Q}}}_{p})$ by sending a character $\nu$ to $(\nu(\gamma_{1}),\cdots,\nu(\gamma_{\deg{\mathfrak{p}}}))\in\mu_{p^{\infty}}^{\deg{\mathfrak{p}}}(\overline{{\mathbb{Q}}}_{p})\subset{\mathbb{G}}_{m}^{\deg{\mathfrak{p}}}(\overline{{\mathbb{Q}}}_{p}).$ A subset $X$ of ${\mathrm{Hom}}(\Gamma_{\mathfrak{p}},\mu_{p^{\infty}})$ is said to be Zariski dense if $X$ is Zariski dense in ${\mathbb{G}}_{m}^{\deg{\mathfrak{p}}}$ over $\overline{{\mathbb{Q}}}_{p}$. subset of characters $\nu$ of $\Gamma_{\mathfrak{p}}$. Our study suggests that a finer non-vanishing holds (cf. Theorem 1.1).
@@
Then there exists an $\ell$-ordinary CM type $\Sigma$ of $K$, i.e.
@@
for $\Sigma_{\ell}:=\{\iota_{\ell}\circ\sigma\ |\ \sigma\in\Sigma\}$ and $\Sigma_{\ell}^{c}:=\{\iota_{\ell}\circ\sigma\circ c\ |\ \sigma\in\Sigma\}$. In the following we first introduce algebraic part $L^{{\mathrm{alg}}}(1,\lambda\nu)\in\overline{{\mathbb{Q}}}$ of the central $L$-value $L(1,\lambda\nu)$.
@@
Our main result towards the problem (Q1) is the following.
@@
Let $K$ be a CM field and $F$ its maximal totally real subfield. Let $\lambda$ be a self-dual Hecke character over $K$ and $\mathfrak{p}$} a prime of $F$ unramified in $K$. Then
@@
Let $\lambda$ be a self-dual Hecke character over a CM field $K$ of infinity type $\Sigma$, which is a CM type of $K$. Let $A_{\lambda}$ be an associated CM abelian variety over $K$. Let $\mathfrak{p}$} be a prime of $F$ split in $K$. Suppose that $\varepsilon(\lambda)=+1$. Then the Mordell–Weil group $A_{\lambda}(K_{{\mathfrak{p}}^{\infty}})$ is finitely generated.
@@
This is a consequence of Theorem 1.1 and [40, Thm. B’]. ∎
@@ 1.2.3. Mod $\ell$ non-vanishing
Our main result towards the problem (Q2) is the following.
@@
Let $K$ be a CM field and $F$ its maximal totally real subfield. Let $\lambda$ be a self-dual Hecke character over $K$ of infinity type $\Sigma+\kappa(1-c)$ for $\Sigma$ a CM type of $K$ and $\kappa\in{\mathbb{Z}}_{\geq 0}[\Sigma]$. Let $\mathfrak{p}$} be a prime of $F$ unramified over $\mathbb{Q}$}. Let $\ell$ be a prime such that $(\ell,6{\mathfrak{p}}D_{F})=1$ and $\Sigma$ is an $\ell$-ordinary CM type. Suppose that
  • $\bullet$
    $\mathfrak{p}$} is either split in $K$ or coprime to ${\rm N}_{K/F}({\mathrm{Cond}}(\lambda))$,
  • $\bullet$
    If $\deg{\mathfrak{p}}>1$, then either ${\mathfrak{p}}{\rm N}_{K/F}({\mathrm{Cond}}(\lambda))$ is divisible by a prime split in $K$ or $\mu_{\ell}(\lambda)=0$.
Then
@@
    If $\deg{\mathfrak{p}}>1$, then Theorems 1.1 and 1.4 are the first results establishing the non-vanishing for all except finitely many $\nu\in\Xi_{\lambda,{\mathfrak{p}}}^{+}$. The prior results [21, 25, 26, 20] established the non-vanishing for a Zariski dense subset of $\Xi_{\lambda,{\mathfrak{p}}}^{+}$, which a priori only implies that the non-vanishing holds for infinitely many $\nu\in\Xi_{\lambda,{\mathfrak{p}}}^{+}$.
@@
    In view of the BSD conjecture (cf. [2]), the above mod $\ell$ non-vanishing would imply that the $\ell$-part of Tate–Shafarevich groups in the $\mathfrak{p}$}-anticyclotomic family is bounded (cf. [6, Cor. 1.2]).
  • iii)
    For ordinary odd primes $\ell$, we expect to remove the hypotheses of Theorem 1.4 via refinement of the strategy elsewhere.
@@ 1.3.1. Iwasawa main conjecture for CM fields
Let $K/F$ be a CM quadratic extension. Let $h_{K}^{-}$ be the relative class number. Let $p$ be an ordinary odd prime.
@@
Based on the work of Hsieh [28], a key arithmetic consequence of Theorem 1.4 to the arithmetic of $K$ is the following divisibility towards the CM Iwasawa main conjecture (cf. [30, p. 90]).
@@
    If $F={\mathbb{Q}}$, then the CM main conjecture is a theorem of Rubin [47], whose proof does not rely on the non-vanishing of Hecke $L$-values.
One has the following application to the BSD conjecture (cf. [28, Cor. 1]).
@@
Theorem 1.6 is asserted in [28]. The approach of loc. cit. is based on the method of Eisenstein congruence on the unitary group $U(2,1)$ over $K$. The first step is construction of a $\Lambda$-adic Eisenstein series ${\mathcal{E}}(\psi)$ on $U(2,1)$ whose constant term is closely related to the Katz $p$-adic $L$-function $L_{\Sigma}(\psi)$. Then a key step is to show $p$-primitivity of the Eisenstein series ${\mathcal{E}}(\psi)$. In loc. cit. it is approached by showing $p$-indivisibility of a certain Fourier–Jacobi coefficient of ${\mathcal{E}}(\psi)$: a toric period of the Fourier–Jacobi coefficient is essentially a product of two Hecke $L$-values. Then the $p$-indivisibility results of [21, 23, 26] are invoked in the case $\deg{\mathfrak{p}}=1$ to show simultaneous $p$-indivisibility of these Hecke $L$-values. Due to the subsequently found gap in Hida’s strategy and consequent weakening [25] of the $p$-indivisibility results of [21, 23, 26] if $\deg{\mathfrak{p}}=1$, Hsieh’s proof of the $p$-primitivity of ${\mathcal{E}}(\psi)$ as in [28, §7.6] is incomplete (see also the introduction of [25]).
@@
1.3.2. Zero divisor of Katz $p$-adic $L$-function
Let $K/F$ be a $p$-ordinary CM quadratic extension.
Let $\lambda$ be a self-dual Hecke character over $K$ of infinity type $\Sigma+\kappa(1-c)$ for $\Sigma$ a $p$-adic CM type of $K$ and $\kappa\in{\mathbb{Z}}_{\geq 0}[\Sigma]$. Let $L_{\Sigma}(\lambda)\in{\mathcal{O}}[\![\Gamma_{K}]\!]$ be the associated Katz $p$-adic $L$-Function, where $\mathcal{O}$} is generated by the image of $\lambda$ over $W(\overline{{\mathbb{F}}}_{p})$ (cf. [34, 29]). Let $\mathfrak{p}$} be a prime of $F$ and
$L_{\Sigma,{\mathfrak{p}}}(\lambda)=\pi_{{\mathfrak{p}}}(L_{\Sigma}(\lambda))$
the associated anticyclotomic $\mathfrak{p}$}-adic $L$-function, where $\pi_{{\mathfrak{p}}}:{\mathcal{O}}[\![\Gamma_{K}]\!]\twoheadrightarrow{\mathcal{O}}[\![\Gamma_{\mathfrak{p}}]\!]$ is the projection map. It interpolates the central $L$-values $L^{\rm alg}(1,\lambda\nu)$ as $\nu$ varies over finite order characters of $\Gamma_{\mathfrak{p}}$.
For a ${\mathbb{Z}}_{p}$-extension $J$ of $K$ contained in the ${\mathbb{Z}}_{p}^{\deg{\mathfrak{p}}}$-anticyclotomic tower $K_{{\mathfrak{p}}^{\infty}}/K$, let $\pi_{J}:{\mathcal{O}}[\![\Gamma_{\mathfrak{p}}]\!]\twoheadrightarrow{\mathcal{O}}[\![{\mathrm{Gal}}(J/K)]\!]$ denote the projection map.
Theorem 1.9.
Let $K/F$ be a $p$-ordinary CM quadratic extension. Let $\lambda$ be a self-dual Hecke character over $K$ of infinity type $\Sigma+\kappa(1-c)$ for $\Sigma$ a $p$-adic CM type of $K$ and $\kappa\in{\mathbb{Z}}_{\geq 0}[\Sigma]$ such that $\varepsilon(\lambda)=+1$. For a prime $\mathfrak{p}$} of $F$, let $L_{\Sigma,{\mathfrak{p}}}(\lambda)$ be the associated anticyclotomic Katz $\mathfrak{p}$}-adic $L$-function. Then for any ${\mathbb{Z}}_{p}$-extension $J$ of $K$ contained in the ${\mathbb{Z}}_{p}^{\deg{\mathfrak{p}}}$-extension $K_{{\mathfrak{p}}^{\infty}}/K$, we have
$\pi_{J}(L_{\Sigma,{\mathfrak{p}}}(\lambda))\neq 0.$
In other words, the $(\deg{\mathfrak{p}})$-variable anticyclotomic $\mathfrak{p}$}-adic $L$-function $L_{\Sigma,{\mathfrak{p}}}(\lambda)$ does not vanish on any ${\mathbb{Z}}_{p}$-line in the tower $K_{{\mathfrak{p}}^{\infty}}/K$. This rigidity for the zero divisor of $L_{\Sigma,{\mathfrak{p}}}(\lambda)$ is an apparent consequence of Theorem 1.1.
Remark 1.10.
In particular, Theorem 1.9 implies the non-vanishing of $L_{\Sigma,{\mathfrak{p}}}(\lambda)$ (see also [4, Thm. 1.2]).

### v2

@@ Abstract.
Let $\lambda$ be a self-dual Hecke character over a CM field $K$. Let $\mathfrak{p}$} be a degree one prime of the maximal totally real subfield $F$ of $K$ and $\Gamma_{{\mathfrak{p}}}$ the Galois group of the anticyclotomic ${\mathbb{Z}}_{p}$-extension of $K$ unramified outside $\mathfrak{p}$}. We prove that
@@
Our approach and results complement the prior work initiated by Hida’s ideas on the arithmetic of Hilbert modular Eisenstein series, studied via mod $\ell$ analogue of the André–Oort conjecture. The previous results established the non-vanishing only for infinitely many characters $\nu$. Our approach is based on the arithmetic of a CM modular form on a Shimura set, studied via arithmetic of the CM field and Ratner’s ergodicity of unipotent flows.
@@ 1. Introduction
A basic example arises from CM motives or Hecke characters over a CM field. The aim of this paper is to establish mod $\ell$ non-vanishing of central Hecke $L$-values in a $p$-adic self-dual family over a CM field for primes $\ell\neq p$ (see Theorems 1.1 and 1.4). It has an application to CM Iwasawa theory: the completion of Hsieh’s proof of Eisenstein congruence divisibility towards the CM main conjecture [27] (see Theorem 1.6).
@@ 1.1. Context
Hida’s mod $\ell$ non-vanishing of Hecke $L$-values over a general CM field has found applications to the arithmetic of the CM field (cf. [21, 39]). It is an ingredient in Hsieh’s 2014 proof of Eisenstein congruence divisibility [27] towards the CM Iwasawa main conjecture over a CM field. However, in 2018 a gap was found in Hida’s strategy, specifically in his proof of Zariski density of CM points on a self-product of $\mathcal{X}$}. In a recent paper [24] Hida recovered [^2] For the case $F={\mathbb{Q}}$, see also [5, 40]. several of the results of [20, 22, 25]. However, if $\deg{\mathfrak{p}}=1$, then the non-vanishing is only proved for infinitely many characters $\nu$ and not all except finitely many $\nu$ as originally claimed. Consequently, Hsieh’s proof [27] of $\ell$-indivisibility of an ancillary Eisenstein series on the unitary group $U(2,1)$, and hence that of the main result of [27] towards the CM Iwasawa main conjecture are incomplete.
In this paper we establish a uniform non-vanishing of self-dual Hecke $L$-values over a general CM field $K$ by a new method. For a self-dual Hecke character $\lambda$ over $K$ and a degree one prime $\mathfrak{p}$} of $F$, we prove that
@@
for all except finitely many finite order characters $\nu$ of $\Gamma_{\mathfrak{p}}$ such that $\varepsilon(\lambda\nu)=+1$ (see Theorem 1.1). We also prove an analogous mod $\ell$ non-vanishing for primes $\ell\neq p$ ordinary with respect to the CM quadratic extension $K/F$ (see Theorem 1.4). The latter leads to completion of Hsieh’s proof of Eisenstein congruence divisibility towards the CM main conjecture (see  Theorem 1.6).
Our approach is rooted in the arithmetic of a CM modular form on Shimura set arising from a totally definite quaternion algebra over $F$ (cf. [5]). It relies on the arithmetic of the CM field, automorphic representation theory, the Waldspurger formula and Ratner’s ergodicity of unipotent flows  [44]. To link arithmetic of the Shimura set with that of the CM field, a key is an $\ell$-integral comparison of quaternionic and CM periods (see Theorem 1.10), which maybe of independent interest.
@@ 1.2.1. Set-up
Let $\lambda$ be a self-dual Hecke over $K$, i.e.
@@
Note that $\lambda\cdot|\ |_{{\mathbb{A}}_{K}}^{1/2}$ is conjugate symplectic self-dual in the sense of [16]. The infinity type of $\lambda$ is of the form $\Sigma+\kappa(1-c)$ for $\Sigma$ a CM type of $K$ and $\kappa=\sum_{\sigma\in\Sigma}\kappa_{\sigma}\cdot\sigma\in{\mathbb{Z}}_{\geq 0}[\Sigma]$, i.e.
@@
$\Xi_{\lambda,{\mathfrak{p}}}^{+}=\{\nu\in\Xi_{\mathfrak{p}}\ |\ \varepsilon(\lambda\nu)=+1\},$
@@
The literature on related problems [20, 22, 23, 24, 11, 12] suggests that such a non-vanishing holds for a Zariski dense [^3] Let $\{\gamma_{1},\cdots,\gamma_{\deg{\mathfrak{p}}}\}$ be a set of topological generators of $\Gamma_{\mathfrak{p}}$. Regard the set of continuous characters ${\mathrm{Hom}}(\Gamma_{\mathfrak{p}},\mu_{p^{\infty}})$ as a subset of ${\mathbb{G}}_{m}^{\deg{\mathfrak{p}}}(\overline{{\mathbb{Q}}}_{p})$ by sending a character $\nu$ to $(\nu(\gamma_{1}),\cdots,\nu(\gamma_{\deg{\mathfrak{p}}}))\in\mu_{p^{\infty}}^{\deg{\mathfrak{p}}}(\overline{{\mathbb{Q}}}_{p})\subset{\mathbb{G}}_{m}^{\deg{\mathfrak{p}}}(\overline{{\mathbb{Q}}}_{p}).$ A subset $X$ of ${\mathrm{Hom}}(\Gamma_{\mathfrak{p}},\mu_{p^{\infty}})$ is said to be Zariski dense if it is so in ${\mathbb{G}}_{m}^{\deg{\mathfrak{p}}}$ over $\overline{{\mathbb{Q}}}_{p}$. subset of characters $\nu$ of $\Gamma_{\mathfrak{p}}$. Our study suggests that a finer non-vanishing holds (cf. Theorem 1.1).
@@
Then there exists an $\ell$-adic CM type $\Sigma$ of $K$, i.e.
@@
for $\Sigma_{\ell}:=\{\iota_{\ell}\circ\sigma\ |\ \sigma\in\Sigma\}$ and $\Sigma_{\ell}^{c}:=\{\iota_{\ell}\circ\sigma\circ c\ |\ \sigma\in\Sigma\}$. In the following we first introduce the algebraic part $L^{{\mathrm{alg}}}(1,\lambda\nu)\in\overline{{\mathbb{Q}}}$ of the central $L$-value $L(1,\lambda\nu)$.
@@
Our main result towards the problem (Q1) is the following (cf. Corollary 4.20).
@@
Let $K$ be a CM field and $F$ its maximal totally real subfield. Let $\lambda$ be a self-dual Hecke character over $K$ and $\mathfrak{p}$} a degree one prime of $F$ unramified in $K$. Then
@@
Let $\lambda$ be a self-dual Hecke character over a CM field $K$ of infinity type $\Sigma$, which is a CM type of $K$. Let $A_{\lambda}$ be an associated CM abelian variety over $K$. Let $\mathfrak{p}$} be a degree one prime of $F$ split in $K$. Suppose that $\varepsilon(\lambda)=+1$. Then the Mordell–Weil group $A_{\lambda}(K_{{\mathfrak{p}}^{\infty}})$ is finitely generated.
@@
This is a consequence of Theorem 1.1 and [38, Thm. B’]. ∎
@@ 1.2.3. Mod $\ell$ non-vanishing
Our main result towards the problem (Q2) is the following (cf. Theorem 5.7).
@@
Let $K$ be a CM field and $F$ its maximal totally real subfield. Let $\lambda$ be a self-dual Hecke character over $K$ of infinity type $\Sigma+\kappa(1-c)$ for $\Sigma$ a CM type of $K$ and $\kappa\in{\mathbb{Z}}_{\geq 0}[\Sigma]$. Let $\mathfrak{p}$} be a degree one prime of $F$ such that $\mathfrak{p}$} is either coprime to ${\rm N}_{K/F}({\mathrm{Cond}}(\lambda))$ or split in $K$. Let $\ell$ be a prime such that $(\ell,6{\mathfrak{p}}D_{F})=1$ and $\Sigma$ is an $\ell$-adic CM type. Then
@@
    In view of the BSD conjecture (cf. [2]), the above mod $\ell$ non-vanishing would imply that the $\ell$-part of Tate–Shafarevich groups in the $\mathfrak{p}$}-anticyclotomic family is bounded (cf. [5, Cor. 1.2]).
@@
    We expect to remove the hypotheses of Theorem 1.4 via refinement of the strategy elsewhere.
@@ 1.3.1. Iwasawa main conjecture for CM fields
Let $K/F$ be a CM quadratic extension and $h_{K}^{-}$ the relative class number. Let $p$ be an ordinary odd prime.
@@
Based on the work of Hsieh [27], a key arithmetic consequence of Theorem 1.4 to the arithmetic of $K$ is the following divisibility towards the CM Iwasawa main conjecture (cf. [29, p. 90]).
@@
    If $F={\mathbb{Q}}$, then the CM main conjecture is a theorem of Rubin [45], whose proof does not rely on the anticyclotomic non-vanishing of Hecke $L$-values.
One has the following application to the BSD conjecture (cf. [27, Cor. 1]).
@@
Theorem 1.6 is asserted in [27]. The approach of loc. cit. is based on the method of Eisenstein congruence on the unitary group $U(2,1)$ over $K$. It begins with construction of a $\Lambda$-adic Eisenstein series ${\mathcal{E}}(\psi)$ on $U(2,1)$ whose constant term is closely related to the Katz $p$-adic $L$-function $L_{\Sigma}(\psi)$. A key is to show $p$-primitivity of the Eisenstein series ${\mathcal{E}}(\psi)$. In loc. cit. it is approached by showing $p$-indivisibility of a certain Fourier–Jacobi coefficient of ${\mathcal{E}}(\psi)$. Since a toric period of the Fourier–Jacobi coefficient is essentially a product of two Hecke $L$-values, the $p$-indivisibility results of [20, 22, 25] are invoked in the case $\deg{\mathfrak{p}}=1$ to show simultaneous $p$-indivisibility of these Hecke $L$-values. Due to the subsequently found gap in Hida’s strategy and consequent weakening [24] of the $p$-indivisibility results of [20, 22, 25] in the case $\deg{\mathfrak{p}}=1$, Hsieh’s proof of the $p$-primitivity of ${\mathcal{E}}(\psi)$ as in [27, §7.6] is incomplete (see also the introduction of [24]).